Shift summary Day November 9, 1999

Nickolas Solomey (solomey@hep.uchicago.edu)
Tue, 9 Nov 1999 15:33:47 -0600

Shift summary Day November 9, 1999

On shift: Roland and Nick

Hours of beam available at all intensities: 8.5 hr
Hours of beam available at nominal intensity: 8.5 hr
Hours of E799 data (special runs not included): 8.0 hr
Hours of special runs (laser scans, muon runs...): 0.0 hr

Accounting of lost time: .5 hr for CsI problems channel 1137, and
a few kfrends.

========================================================================
Runs ended this shift: none

Run 14933 running during shift

Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift: Data taking, the accelertor never missed a spill
and the intensity was always at or above 9e12 (I am impressed, ns).

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14933 07:00 New shift crew takes over
156 07:58 The problem CsI crystal 1096, seems to be
flat gain for the last half hour, but we
will continue to monitor.
209 09:09 Ped problem and laser rms to wide.
09:10 Paged Julie, who phoned quickly had us
reset pipline, kfrend and hv check on
the bad channel 1137.
218 09:23 Running again 1137 seems to be ok now.
238 HCC inefficient, but it went away the next
spill.
261 Ped Problem again, csi channel 1137 is
like before 97% out of range. Repeated
recovery procedure of last time, reset
pipeline with dbc and kfrend. Then it was
fine.
264 Good data taking.
279 Trigger in error state, did a kfrend and
it was fine afterwards.
281 Dido previous comment for spill 279.
319 Triggers stopped dyc stuck.
328 Trigger problems, traced to Ashkan having
started on ktev1 a job for new xt and
his job was running on processor zero
causing us to miss poritions of the
spill.
333 Data taking is fine now that Ashkan killed
his xt job.
351 12:30 This spill had a bad structure.
12:40 Beam spill structure getting worse, called
MCR and asked them to look into it.
419 Trigger in error state, did a kfrend.
430 14:08 Spill structure getting worse so called MCR
they said some of their people are testing
something and it might be causing it, they
would not elaborate.
15:00 Beam quality is much improved lately, but
not as good as the first 5 hours of this
shift.
492 15:30 End of shift, all still fine.

------------------------------------------------------------------------
Concern from previous shift is to watch the slow drift of CsI channel
1096, it is out of spec by more than 10% but JW said it is alright as
long as it does not jump.

------------------------------------------------------------------------
Follow up on next shift. If accelertor goes off or low intensity
for a while then phone Ashkan so he can run his xt job.

Dave N. showed up and said that the Virgina crew might want to
check on the Lq N2, Dave asked us to leave them a note about it.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey